### v1

# Automating the Diagnosis of Human Vision Disorders by Cross-modal 3D Generation

Abstract
Understanding the hidden mechanisms behind human’s visual perception is a fundamental quest in neuroscience, underpins a wide variety of critical applications, e.g. clinical diagnosis. To that end, investigating into the neural responses of human mind activities, such as functional Magnetic Resonance Imaging (fMRI), has been a significant research vehicle. However, analyzing fMRI signals is challenging, costly, daunting, and demanding for professional training. Despite remarkable progress in artificial intelligence (AI) based fMRI analysis, existing solutions are limited and far away from being clinically meaningful. In this context, we leap forward to demonstrate how AI can go beyond the current state of the art by decoding fMRI into visually plausible 3D visuals, enabling automatic clinical analysis of fMRI data, even without healthcare professionals. Innovationally, we reformulate the task of analyzing fMRI data as a conditional 3D scene reconstruction problem. We design a novel cross-modal 3D scene representation learning method, Brain3D, that takes as input the fMRI data of a subject who was presented with a 2D object image, and yields as output the corresponding 3D object visuals. Importantly, we show that in simulated scenarios our AI agent captures the distinct functionalities of each region of human vision system as well as their intricate interplay relationships, aligning remarkably with the established discoveries of neuroscience. Non-expert diagnosis indicate that Brain3D can successfully identify the disordered brain regions, such as V1, V2, V3, V4, and the medial temporal lobe (MTL) within the human visual system. We also present results in cross-modal 3D visual construction setting, showcasing the perception quality of our 3D scene generation.
1 Introduction
Understanding how humans see and perceive the world is a fundamental question in neuroscience and vision [1, 2]. The ability to perceive and interpret the environment in three dimensions (3D) is a critical aspect of human vision that enables human to navigate, interact, and make sense of the world around us [4, 3, 5]. Despite significant advancements in the understanding of the neural mechanisms underlying visual perception [6, 7, 8, 9], the intrinsic processes involved in 3D vision remain largely unexplored.
Traditional approaches in neuroscience and life sciences, such as animal studies [10, 11, 12] and invasive techniques [13, 14], have provided valuable insights but are limited by several factors. These limitations include the need for animal sacrifice [15], ethical concerns [16], and interspecies heterogeneity [17], which can hinder the translation of findings to human vision. Additionally, traditional methods often lack the spatial and temporal resolution necessary to capture the complex dynamics of neural activity underlying 3D vision perception. These limitations have restricted the progress of research in this field and highlight the need for novel interdisciplinary approaches.
Artificial vision, which aims to develop computational models inspired by biological vision [18, 19, 20], has emerged as a promising avenue for understanding the neural basis of visual perception. While artificial vision models have achieved remarkable success in various computer vision tasks, such as object recognition [21, 22, 23], detection [24, 25, 26], and scene understanding [27, 28], their contribution to understanding biological vision remains limited. These models primarily focus on replicating human-like performance on specific tasks using visual input, without directly incorporating brain signals or mimicking the neural processes involved in biological vision. Furthermore, these models are not designed to simulate the effects of brain signal perturbations on vision, which is crucial for diagnosing and understanding vision-related brain disorders.
Figure 1: Brain3D: Processing brain fMRI for clinical diagnosis. The subject’s visual system first precepts the object stimulus, triggering specific patterns of activity in the brain’s visual processing areas which is then captured and expressed by fMRI data. Brain3D then leverages a fMRI encoder to extract the feature, feeds it into a 2D diffusion model for detailed 3D generation. By contrasting the object stimulus and generated 3D visuals, even a non-expert can conduct precise diagnosis of atypical neurological conditions.
To bridge this gap and gain deeper insights into how the human brain processes and interprets visual information, here we explore the intrinsic 3D human vision perception through the lens of fMRI data [29], known for its high spatial resolution, relatively low cost compared to other neuroimaging techniques, and non-invasiveness. We introduce Brain3D, a bio-inspired method capable of understanding how the brain interprets and processes visual stimuli in 3D. Brain3D takes fMRI data of a subject presented with a 2D object image and yields the corresponding 3D object model. Importantly, our AI agent captures the distinct functionalities and interplay relationships of each region in the human vision system, aligning with established neuroscience discoveries. This alignment allows Brain3D to serve as a powerful tool for clinical fMRI diagnosis, offering innovative diagnostic and therapeutic approaches in neurology and psychology. Preliminary experiments demonstrate that our AI agent can effectively diagnose regions associated with disorders, including V1, V2, V3, V4, and the medial temporal lobe (MTL). This research pioneers the understanding of the neural basis of human 3D vision with artificial neural networks, shedding light on cognitive processes and paving the way for theoretical exploration and practical applications in neuroscience.
Figure 2: Overview of our Brain3D that decodes the functional MRI (fMRI) signals into 3D reconstructions. Initially, the subject’s visual system precepts the object stimulus, triggering specific patterns of activity in the brain’s visual processing areas. These patterns are recorded as fMRI signals. To generate the 3D object model, Brain3D leverages two specialized encoders: the high-level encoder captures abstract high-level visual concepts, while the low-level encoder focuses on processing low-level visual details. The extracted features are passed into a two-stage reconstruction pipeline to generate the 3D representation of the object. The training procedure involves optimizing low-level and high-level encoders with the feature distillation objective $L_{image}$ and $L_{text}$ respectively, each incorporating bidirectional CLIP [72] contrastive learning and MixUp [71] augmentation for enhanced generalization.
Stimulus
MindEye [35]
Brain3D
Stimulus
MindEye [35]
Brain3D
Stimulus
MindEye [35]
Brain3D
Figure 3: Brain3D generates finer 3D visuals from fMRI under 2D stimuli compared with previous methods. For each case, the first column displays the 2D stimuli presented to the subjects. The second column presents 2D reconstruction from MindEye [35]. The three and fourth column exhibits two distinct views of our 3D visuals generated from fMRI, followed by the final column depicting the surface normals of our generation, yielding higher quality output than MindEye [35]. The zoom-in visualizations in the last two rows demonstrates our better quality and more accurate fMRI-based generation.
2 Results
Overview of fMRI-based diagnosis through Brain3D.
As a bio-inspired method our Brain3D can understand how the brain interprets and processes visual stimuli in 3D, thus allowing clinical diagnosis (Fig. 2). It efficiently maps 2D visual stimuli into 3D generations by analyzing the intricate functionalities of distinct brain regions.
Specifically, our model first deploys two specialized neural encoders to encode fMRI data into multi-level neural embeddings. The low-level encoder is designed to encode fMRI data from the basic visual cortex (V1 to V4), focusing on visual details such as color, shape, and texture, into low-level neural embeddings. Concurrently, the high-level encoder concentrates on the medial temporal lobe (MTL) to process visual information, such as object semantics and types. It encodes fMRI data from the MTL regions to high-level embeddings. Additionally, we introduce a UMAP-based projection [65] to further reduce noise in the high-level embeddings by decomposing these embeddings into separate distinct representations.
Subsequently, we design a two-stage diffusion method [30, 31, 32] to reconstruct the corresponding 3D objects. 3D objects are generated through distilling a 2D diffusion model with condition on both the high level and low level fMRI embeddings, using Score Distillation Sampling (SDS) [33]. For high-fidelity visual production, we employ a coarse-to-fine optimization strategy, with the Neural Radiance Field (NeRF) [34] as 3D scene representation for initial geometry learning followed by 3D mesh representation learning for detail recovery.
With generated 3D visuals, we conduct clinical fMRI diagnosis on atypical neurological conditions. We randomly select samples from NSD testing set with simulated disorder over five brain regions: V1, V2, V3, V4, and MTL. Then, we invite 100 non-experts to detect which brain region has disorder by comparing the original image and the generated multi-view images by our model. The results affirm Brain3D’s effectiveness in identifying and diagnosing neurological abnormalities across multiple brain regions, including V1, V2, V3, V4, and the medial temporal lobe (MTL).
Demanding challenges with fMRI-based 3D generation.
Generating 3D visual models from brain signals poses formidable challenges, especially concerning the collection and interpretation of fMRI signals. Recording brain responses from multiple angles simultaneously is technically demanding and costly, particularly in scenarios resembling real-world environments where only limited viewpoints are available. Additionally, the intricate distribution of brain regions further complicates the processing, resulting in a significant diversity of high-level information. Decoding and effectively utilizing this variety of information becomes exceedingly challenging, particularly when the amount of fMRI data is limited and costly to acquire. These challenges underscore the substantial difficulties in advancing fMRI-based 3D modeling techniques within the field of neuroscience.
Figure 4: Collaborative property of the left and right hemispheres with Brain3D. (a) To examine the difference and interaction between the human brain’s two hemispheres, we conduct specialized experiments assessing 3D generation quality using fMRI from either hemisphere or both. (b) The function of the left and right hemispheres varies significantly for different specific objects under various metrics. In general, the left favors finer details and intricate structures, while the right focuses more on overall shape and silhouette. (c) Notably, using a combination of data from both hemispheres led to an enhanced generation. This synergistic effect was evident in improved performance across various metrics. More examples can be found in Fig. s7.
Brain3D decode 3D models solely from fMRI signals with finer semantics in comparison with previous 2D models.
To test the effectiveness of Brain3D in 3D generation, we perform comprehensive comparisons with an existing 2D method [35]. We establish 3D-based evaluation benchmarks, including 3D Inception [36], 3D CLIP [37], 3D EfficientNet [38], and 3D Swav [39], for subsequent evaluations (see Sec. 4.6 for detailed descriptions of these benchmarks).
From the fMRI data Brain3D extracts rich appearance, semantic and geometric information, presenting a high degree of consistency with the stimuli images (Fig. 3). Despite limited observation with the giraffe, skier and teddy bear in the three cases (see the last two rows of Fig. 3), our bio-inspired model can process 3D details and comprehend their semantics beyond what is immediately visible in the 2D image, achieving finer semantic details than MindEye [35]. This result underscores the remarkable human’s neural processing capability in constructing intricate 3D structures by interpreting and internalizing 3D geometric information from 2D visual stimuli [40].
Enhanced visual quality through collaboration between the left and right hemispheres.
Figure 5: Visual region predominates in comprehension. (a) Overall, visual regions outperform the medial temporal lobe (MTL) regions under varying metrics. Further, V1 contributes the most among visual regions. (b) The vision region demonstrates a higher 3D generation quality with elevated performance. The significance level for MTL vs vision is: 3D Inception: p=1e-23, 3D CLIP: p=1e-16, 3D EfficientNet p=1e-22, 3D Swav p=3e-15. Performance of each generated object in these criteria is represented by crosses (+). (c) V1 Regions exhibit better performance than V2, V3 and V4 in generation. The significance level is: 3D Inception V1 vs. (V2, V3, V4) p=(1e-17, 4e-9,9e-5), (V1, V2, V3, V4) vs. all p=(3e-8,1e-47,4e-32,6e-24). 3D CLIP V1 vs. (V2, V3, V4) p=(1e-30,6e-23,2e-18), (V1, V2, V3, V4) vs. all p=(9e-3,3e-31,2e-23,7e-19). 3D EfficientNet V1 vs. (V2, V3, V4) p=(6e-23,1e-18,7e-10), (V1, V2, V3, V4) vs. all p=(5e-4,1e-51,1e-33,2e-21). 3D Swav V1 vs. (V2, V3, V4) p=(2e-11,2e-6,1e-4), (V1, V2, V3, V4) vs. all p=(1e-4,4e-26, 5e-18, 8e-15). Significance levels are indicated as * = p<1e-2, ** = p<1e-3, *** = p<1e-4. (d) MTL region tends to extract semantic visual concepts, while the visual cortex concentrates on textures and silhouettes. (e) V1 dominates in inferring the object silhouette, whilst V2, V3, and V4 are more focused on detailed aspects of texture. More examples can be found in Fig. s8.
We investigate whether our model presents biological consistency in terms of regional functionalities. We first examine the roles of the left and right brain hemispheres, as well as their collaboration during the 3D object reconstruction process (Fig. 4a).
Specifically, we observe that the two hemispheres make different holistic contributions to object perception, due to exhibiting distinct performances for specific objects (Fig. 4b). The left hemisphere tends to depict finer details and intricate structures in object visualization, while the right hemisphere better captures the overall shape and silhouette. These results align with previous discoveries that the left hemisphere is involved in detail-oriented tasks while the right hemisphere in more holistic tasks [43, 44]. However, our results challenge the notion of strict functional division. We found notable overlaps and interactions between the two hemispheres. Remarkably, numerous scenes rendered by the right hemisphere, as illustrated in Fig. 4a and Fig. s7, also exhibit intricate textures—typically a characteristic attributed to the left hemisphere. Conversely, the left hemisphere demonstrated capability in detecting overarching shapes, which is usually a strength of the right hemisphere.
Despite these different functional tendencies, our quantitative analysis, using metrics like 3D Inception, 3D CLIP, 3D EfficientNet, and 3D Swav in Fig. 4c, shows that both hemispheres contribute similarly to 3D object perception. Their collaboration significantly contributes to the quality of 3D objects (Fig. 4c). This result echos the biological discovery on the interaction between the left and right brain [41, 42, 43].
Figure 6: Illustration of Brain3D to process fMRI for clinical diagnosis. (a) In non-expert diagnosis, participants are first equipped with knowledge about the function with varying visual system regions (V1, V2, V3, V4, MTL) and a number of case study for their diagnosis practice. We show cases both with and without disorders for information. (b) Regional disorder diagnosis performance by non-professionals using our Brain3D. (c) 3D representation of Brain3D achieves improved F1 diagnosis score than 2D model. (d) Visual example of disorder diagnosis with Brain3D.
Dominance of V1 visual regions in feature and silhouette processing.
We further investigate the functionalities of the various visual regions and the medial temporal lobe (MTL) region (Fig. 5a).
Previous studies have demonstrated that the visual cortex [45] is predominantly involved in processing color and texture information, whereas the medial temporal lobe (MTL) region [46, 47, 48] is more focused on semantic information. In comparison, our research shows that the visual cortex exhibits higher visual decoding ability than MTL region (Fig. 5b). The significance level refers to the probability of error when rejecting the null hyothesis that visual cortex shares the similar performance as MTL region. Specifically, the visual cortex captures local fine-grained visual information such as the “windows” of an airplane and the “motorcycle rider”, while the MTL region captures holistic shapes like “airplane skeleton” and “person” (Fig. 5d). This visualization highlights higher-level 3D-shape-related concepts while foregoing pixel-level details. This finding provides a novel perspective on how the MTL region may contribute to the formation and recall of long-term memories, emphasizing its role in abstracting complex visual information.
In line with earlier researches [6, 7], V1 region exhibits predominant ability in visual representation (Fig. 5c) among various visual regions. To further break down, V1 region plays a pivotal role in handling elementary visual information, including edges, features, and color. V2 and V3 regions play a role in synthesizing various features into complex visual forms, while V4 region focuses on basic color process at coarse level (Fig. 5e). However, our research brings to light a nuanced understanding of the interdependence within the visual cortex. Specifically, we demonstrate the reliance of the V2, V3, and V4 regions on the V1 region for constructing 3D vision. This is evident from the disrupted 3D geometries observed in Fig. 5e. This reliance likely stems from the fact that these regions process information initially received by V1 [1], indicating a foundational dependence on V1 for effective visual processing. These findings emphasize the collaborative and interactive roles of these brain regions in forming a comprehensive visual representation. This result is thus biologically consistent to great extent.
Brain3D can diagnose regions with disorder (preliminary study).
Motivated by the high biological consistency in terms of regional functions, we examine the usefulness of Brain3D in diagnosing disordered brain regions by designing a series of simulation based experiments (Fig. 6a). To compare performance of Brain3D and 2D models, we designed three branches of experiment: 2D model MindEye [35], Brain3D front view and Brain3D back view. In each branch, the participants were tasked with determining the impacted brain region of the patient from V1, V2, V3, V4, and MTL by observing 3D visuals generated from their fMRI signals. Their average correctness of judgement is represented as evaluation of clinical diagnosis. Notably, almost perfect diagnosis is achieved for the disorders with the V1 and MTL regions (Fig. 6b). However, distinguishing the disorders among V2, V3, and V4 regions presents more challenges, as reflected by the incidence of incorrect diagnoses in these areas. This difficulty is attributed to the significant overlap and interaction in their processing roles, as they all engage in visual information processing following the V1 region. Nonetheless, when a disorder was localized to a single region among V2, V3, or V4, participants generally succeeded in identifying that the issue stemmed from one of these three regions, despite the nuanced differences in their functionalities. Diagnosis among all regions in Brain3D outperforms previous 2D methods (Fig. 6c). We provide visual examples for region disorder diagnosis (Fig. 6d). V1 disorder tends to cause significant distortion to the output, consistent with the earlier finding about its critical role in processing basic visual information. Disorder in V2 region leads to changes in the detailed shape and motion of objects, e.g., orientation of the airplane and shape of the motorcycle. V3 region primarily affects the details and movements, e.g., the change from motorcycle to rider. Disorder in V4 region might cause color shift and variation. Instead, MTL’s disorder could be concerned with high-level semantics, e.g., increased ambiguity in object appearance. See Sec. 4.8 for more details.
Brain3D’s ability to generate 3D objects from fMRI data significantly contributes to this diagnostic efficiency. This capability is especially beneficial in assessing the MTL region. Without 3D information from the front view, disorders in the MTL region exhibit minimal impact, as the ‘Front’ column in Fig. s1c remains nearly unchanged. This is likely because the visual region still efficiently extracts the silhouette from the front perspective. However, the absence of high-level information due to a disordered MTL region leads to a much vaguer back view, creating a stark contrast to the normal state. The inclusion of this back view in diagnosis significantly improves accuracy, as demonstrated in Fig. s1a with a limited number of images. The use of the back view in diagnostic assessments also enhances the accuracy of diagnoses across all visualizations from patients, as shown in Fig. s1b. This improved diagnostic capability might stem from the impaired MTL region’s reduced ability to use semantic information for understanding unseen perspectives in a given 2D image. In 2D model [35], although abnormalities in the MTL brain area also lead to changes in the generated 2D images (Fig. s1c), this changes do not have obvious features, making it difficult to distinguish and diagnose from other brain area abnormalities (such as V1 abnormalities). This directly leads to a significant decrease in the success rate of diagnosing the MTL brain region (Fig. s1a). In comparison, Brain3D reasonably exploits MTL region, thus being able to distinguish disorders in MTL from other brain regions more directly by the blurring in the back view (Fig. s1c). This demonstrates that the 3D visualization capabilities of Brain3D are vital for clinical diagnosis and comprehending 3D human vision.
To explore more cost-effective diagnostic methods, our study extended to experiments where participants were provided with a limited number of visualizations for brain disorder diagnosis. This approach was driven by the desire to ascertain if accurate diagnostics could be attained with fewer resources such as time and cost. The results were significant, as demonstrated in Fig. s2: even with a reduced number of images, participants were able to accurately tell disorders in the V1 and MTL regions, significantly cutting down the time and resources usually needed for thorough diagnoses. Specifically, even with as few as 10 images (Fig. s2a), participants achieved over 95% of diagnostic accuracy for disorders in the V1 and MTL regions. This implies that by utilizing merely 10% of the potential images from patients, we can still reach a plausible diagnisis of V1 and MTL region thereby saving almost 90% of the time and effort as required in our full diagnostic process.
Ablation analysis.
Our perceptual pipeline employs Score Distillation Sampling (SDS) [33] with the inclusion of both low-level and high-level guidance. The integration of both low-level and high-level guidance achieves the best performance across all categories (Fig. s3a). As shown in Fig. s3b, when solely relying on high-level guidance, the representations of the train, airplane, and clock tower appear significantly blurred, displaying only coarse shapes. These coarse shapes stem from the geometric and semantic information embedded within the high-level representation. Low-level guidance primarily captures the color and texture attributes of the objects. Consequently, the outcomes with low-level guidance (see the first column), featuring the front view, align well with the provided image, while other views fall short in terms of shape and semantic fidelity. The combination of both high-level and low-level guidance results in significantly improved representations of the airplane, train, and clock tower, where both the semantic and textural structures closely resemble those of the testing image.
The Semantic Stage within our framework is designed to extract finer details from fMRI embeddings, enhancing its contribution to 3D generation. After the Semantic Stage the detailed textures of the building in the first row and the outline of the steamship in the second row in Fig. s4c exhibit much greater clarity. These details from fMRI are challenging to represent in a coarse NeRF scene but can be effectively distilled within a 3D mesh.
[…]
3 Discussion
Understanding how human vision system extracts and processes the physical world is a critical question in neuroscience, with vital applications in healthcare and other fields. For the first time, this research models the underlying mechanisms of human visual perception by mapping fMRI of human brain activities to the corresponding 3D objects. This goes beyond previous studies that are limited to statistical fMRI analysis [42, 1, 2, 7].
Despite the advancements in our 3D scene generation and clinical diagnosis, discrepancy remains between the generated objects and the actual images viewed by participants. This phenomenon can be attributed in part to cognitive biases [78] and the inherent blurring effects [79] that occur during the image-to-fMRI cognitive stage. An example of this can be seen in the last row of Fig. 3 where the bird’s beak appears shut. This is because the upper half of the beak appears similar as the background, illustrating how visual processing and brain interpretation can influence our visual perception. Signal resolution also plays a significant role. The 1.8 mm resolution of fMRI data leads to a loss of fine details, particularly in smaller neuronal structures. In addition, current 3D generation models also struggle to accurately capture background details that extend beyond the camera’s field of view. These limitations highlight the challenges in accurately representing and analyzing complex visual information.
By enabling deeper and more direct analysis of visual perception, our study opens up a more advanced and convenient approach to the clinical diagnosis of brain disorders. Previous fMRI-based diagnostic techniques primarily focused on detecting variations in signal frequency within specific brain regions, such as the motor area for Parkinson’s disease [50, 49], the premotor and motor cortices in cases of Dystonia [51, 52, 53], and motor networks for tremor [54, 55] and Huntington’s disease [56, 57, 58]. However, such methods, relying solely on activation frequency, can be limited due to their narrow focus on direct statistical analysis of fMRI signals, which may not fully account for the variability inherent in fMRI, particularly in the context of disorders. For instance, Parkinson’s disease may exhibit both decreased [59, 60]and increased  [61, 62] in the motor area, while Dystonia also shows conflicting fMRI activity levels [51, 52, 53]. In contrast, Brain3D represents a more sophistic and intuitive diagnostic approach by attempting to recover the target objects from fMRI. The key is our capacity in encoding the functionality of various brain regions that can be easily visualized in the output. As a result, the clinical diagnose process is made significantly simpler and efficient.
Our research also highlights the human brain’s remarkable capability in perceiving 3D-related high-level information from 2D visual stimuli. This perception is facilitated by the synergistic functions of both brain hemispheres and the intrinsic collaboration between different functional regions. For example, the V1 region is fundamentally pivotal along with V2, V3, and V4 regions enhancing the perception through detail and color information, whilst the MTL region focuses on higher-level semantics. For the first time, we have attempted to simulate this intricate mechanisms of human visual perception by innovating a machine learning approach with great potentials for real-world applications.
4 Method
4.1 Pipeline overview
We aim to model the intricate mechanisms behind human vision perception. To achieve this, we develop Brain3D, a model capable of creating 3D representations from the fMRI signal of individual’s minds. Given the fMRI signal collected during an individual viewing a 2D image, we decompose it into two parts according to the sources: visual cortex and MTL. The visual cortex focuses on the low-level texture feature, while the MTL captures the high-level semantics information. For each source, we design a specialized feature embedding encoder.
Formally, this process begins with an fMRI voxel $\mathbf{X}_{i}$. We employ Multilayer Perceptrons (MLPs) as encoders to transform $\mathbf{X}_{i}$ into two types of embeddings: a low-level embedding $f_{L}(\mathbf{X}_{i})$, and a high-level embedding $f_{H}(\mathbf{X}_{i})$. These embeddings are then supervised by respective features under image and textural feature distillation. This multi-level structure thus separates the high-level semantics information and the low-level texture details of the fMRI signal, simulating the biological information processing [64, 63]. The final 3D objects are reconstructed using a two-stage generative diffusion pipeline conditioned on both the low-level and high-level fMRI embeddings.
4.2 Preliminary: Diffusion based 3D generation
Our approach employs a diffusion model [30] aimed at manipulating data through a two-phased process. Initially, a forward process introduces Gaussian noise with parameter $(\alpha_{t},\sigma_{t})$ into the data $\mathbf{x}_{0}$ to achieve noised input $\mathbf{z}_{t}$, which can be mathematically represented as follows:
[…]
Following this, a reverse process is employed to denoise the data, starting from pure noise. This reverse mechanism is driven by a parameterized noise prediction network $\epsilon_{\phi}(\mathbf{x}_{t},t)$ , optimized by minimizing the objective loss function as:
$L_{\rm{Diff}}(\phi,\mathbf{x})=\mathbb{E}_{t\sim U(0,1),\epsilon\sim N(0,1)}\left[\omega(t)|\epsilon_{\phi}(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,t)-\epsilon_{t}|_{2}^{2}\right].$ (2)
This mechanism is further extended to 3D generation through the integration of Score Distillation Sampling (SDS) [33]. It involves two key components: a scene model and a pre-trained Text-to-Image (T2I) diffusion model. The scene model, parameterized by $\theta$, generates differentiable images through $\mathbf{X_{i}}=g(\theta)$. The pre-trained T2I diffusion model predicts the sampled noise $\epsilon_{\phi}(\mathbf{X_{i}}_{t};t,\mathbf{y})$ conditioned on the noisy image $\mathbf{X_{i}}_{t}$, created by adding Gaussian noise with a random scale factor $t$ to $\mathbf{X_{i}}$. The optimization of this extended model is guided by the following objective:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{SDS}(\phi,\mathbf{X_{i}})=\mathbb{E}_{t,\epsilon}[\omega(t)(\hat{\epsilon}_{\phi}(\mathbf{X_{i}}_{t};t,\mathbf{y})-\epsilon_{t})\frac{\partial\mathbf{X_{i}}}{\partial\theta}]\\ &=E_{t,\mathbf{z}_{t}|\mathbf{X_{i}}}[\omega(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}KL(q(\mathbf{z}_{t}|\mathbf{X_{i}}=g(\theta)||p_{\phi}(\mathbf{z}_{t}|\mathbf{y})))].\end{split}$ (3)
Through this computation, our model seeks to align the generated distribution $q(\mathbf{z}_{t}|\mathbf{X_{i}}=g(\theta))$ closely with the target distribution $p_{\phi}(\mathbf{z}_{t}|\mathbf{y})$, ensuring consistency across various camera poses, as detailed in the following integration:
[…]
This integration necessitates consistency in $p_{\phi}(\mathbf{z}_{t}|c,\mathbf{y})$ from each camera pose $c$.
4.3 Tackling high noises with high-level semantics
In modeling human mind signals, a major challenge arises from the intrinsically significant levels of noise and diversity with the high-level information within the human brain (see Fig. s6). This can cause instability in the conditional probability distribution $p_{\phi}(\mathbf{z}_{t}|c,\mathbf{y})$ (see Sec. 4.2 for more details), particularly when considering different camera angles, denoted as $c$. This instability could result in a lack of consistency for 3D object generation. Addressing this issue is critical for ensuring that our model reliably interprets and reconstructs the complex fMRI scans.
We develop a novel approach to solving the above problem using the Uniform Manifold Approximation and Projection (UMAP) [65]. The key idea is to distill the high-level semantic embeddings into a series of representative embedding vectors with a unique projection weight.
Formally, it projects the high-level embedding $f_{H}(\mathbf{X}_{i})$ into a set of stable embeddings, $\left\{f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, derived from the Generative Text-to-Image Transformer [66]. In our design, we concatenate the embeddings extracted from all training images. This results in a comprehensive set of embeddings, $\left\{f_{H}(\mathbf{X}_{i}),f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, for each image. These combined embeddings form a high-dimensional dataset, denoted as D.
Then, we optimize the stable representation of these embeddings, with the goal to minimize the cross-entropy between the high-dimensional probability distributions and their corresponding stable embeddings. It is conducted via stochastic gradient descent for handling the complexity of such high-dimensional data.
For fMRI inference $\mathbf{X}_{i}$, we implement a pre-trained UMAP model, $g_{U}$, to project the high-level embeddings $\left\{f_{H}(\mathbf{X}_{i}),f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$ into a stable space. The resultant embeddings are represented as $\left\{\mathbf{U}_{H},\mathbf{U}_{H_{1}},\mathbf{U}_{H_{2}},...,\mathbf{U}_{H_{n}}\right\}$ with each obtained as:
$\mathbf{U}_{H_{j}}=g_{U}(f_{H_{j}}(\mathbf{X}_{i})).$ (5)
To quantify the similarity between $\mathbf{U}_{H}$ and each $\mathbf{U}_{H_{j}}$, we calculate the cosine similarity, $\rm{s_{i}}$, as follows:
$\rm{s_{i}}=\frac{{\mathbf{U}_{H}\cdot\mathbf{U}_{H_{j}}}}{{|\mathbf{U}_{H}|\cdot|\mathbf{U}_{H_{j}}|}},$ (6)
This similarity measurement is then normalized to establish a sampling probability for each embedding:
$\rm{v}_{j}=\frac{\rm{s}_{j}}{|\mathbf{U}_{H_{j}}|},$ (7)
$\rm{w}_{j}=\frac{\rm{v}_{j}}{\sum_{j=1}^{n}\rm{v}_{j}},$ (8)
This normalization process is critical as it determines how embeddings are sampled during the denoising step of the 2D diffusion model. Under specific conditions, the linear space formed by $\left\{\mathbf{U}_{H_{1}},\mathbf{U}_{H_{2}},...,\mathbf{U}_{H_{n}}\right\}$ can effectively represent $\mathbf{U}_{H}$ linearly, as described in the follows:
$\mathbf{U}_{H}=\sum_{j=1}^{n}\rm{proj}_{\mathbf{U}_{H_{j}}}(\mathbf{U}_{H})=\sum_{j=1}^{n}\frac{\mathbf{U}_{H_{j}}\cdot\mathbf{U}_{H}}{|\mathbf{U}_{H_{j}}|^{2}}\mathbf{U}_{H_{j}}=|\mathbf{U}_{H}|\sum_{j=1}^{n}t_{j}\cdot\mathbf{U}_{H_{j}}.$ (9)
In essence, the UMAP projection serves as a non-linear expression of the high-level embedding $f_{H}(\mathbf{X}_{i})$, contributing significantly to our model’s ability to handle complex data representations.
[…]
$\begin{split}&\nabla_{\theta}\mathcal{L}_{SDS}(\phi,\mathbf{X}_{i})=\mathbb{E}_{t,i,\epsilon}[\omega(t)(\hat{\epsilon}_{\phi}(\mathbf{X}_{i_{t}};t,\mathbf{y_{i}})-\epsilon_{t})\frac{\partial\mathbf{X}_{i}}{\partial\theta}]\\ &=E_{t,i,\mathbf{z}_{t}|\mathbf{X}_{i}}[\omega(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}KL(q(\mathbf{z}_{t}|\mathbf{X}_{i}=g(\theta)||p_{\phi}(\mathbf{z}_{t}|\mathbf{y}_{i})))],\end{split}$ (10)
Here, the expectation of the sampled embedding $i$ is integrated into the computation. This means we selectively utilize these weights to pick the most relevant high-level embeddings. This integration is crucial for ensuring that our model accurately captures and reconstructs the complex patterns inherent in fMRI data, producing more reliable and consistent multi-view visual results.
4.4 Two-stage 3D generation
To effectively leverage the projected high-level embeddings, we devise a two-stage 3D generation architecture characterized by progressively integrating high-level information during 3D object modeling. The first stage is a perceptual stage, utilizing the Neural Radiance Fields (NeRF) [34] for 3D representation. It defines an implicit 3D representation framework expressed as:
[…]
$T(t)=\rm{exp}(-\int_{t_{n}}^{t}\sigma(\textbf{r}(s))ds).$ (12)
This framework allows the gradients from the Score Distillation Sampling (SDS) [33] to directly influence the output $C(\mathbf{r})$, enabling simultaneous optimization of the density Multilayer Perceptron (MLP) $\sigma(\mathbf{r}(t))$ and color MLP $(\mathbf{r}(t),\mathbf{d})$. As a result, NeRF is particularly effective at modifying both the color and shape of the 3D scene in this stage. In our implementation, we incorporate Zero-1-to-3 [67] for low-level guidance and Stable Diffusion [31] for high-level guidance.
Subsequently, a semantic stage follows where we employ DMTet [68] to convert the above NeRF into a 3D mesh [69]. This conversion prepares the 3D model for further processing. We continue to harness Stable Diffusion [31], coupled with a depth-conditioned Control-Net [70] for high-level guidance. While the 3D mesh does not possess the same versatility as NeRF in shaping modifications, it excels in capturing detailed RGB color information. Therefore, after establishing the coarse shape in the perceptual stage, we shift our focus to refining the color details in this stage. This two-tier design allows for an intricate and nuanced development of the 3D scene, leading to a final output that is both elaborately detailed and of high resolution.
4.5 Training and inference
During training, we optimize the parameters for both low-level encoder and high-level encoder, leaving all the rest components frozen. We use a training dataset with image-text pairs. Inspired by MindEye [35], the loss function for low-level $L_{image}$ and high-level $L_{text}$ embeddings are designed with integration of bidirectional CLIP [72] contrastive learning and MixUp [71] augmentation to improve generalization in the beginning of training stage, as described in the follows:
$\begin{split}\mathcal{L}_{mix}=&-\sum_{i=1}^{N}[\lambda_{i}\cdot\log(\frac{\exp(\frac{p_{i}^{*}\cdot t_{m}}{\tau})}{\sum_{m=1}^{N}\exp(\frac{p_{i}^{*}\cdot t_{i}}{\tau})})\\ &+(1-\lambda_{i})\cdot\log(\frac{\exp(\frac{p_{i}^{*}\cdot t_{k_{i}}}{\tau})}{\sum_{m=1}^{N}\exp(\frac{p_{i}^{*}\cdot t_{m}}{\tau})})]\\ &-\sum_{j=1}^{N}[\lambda_{j}\cdot\log(\frac{\exp(\frac{p_{j}^{*}\cdot t_{j}}{\tau})}{\sum_{j=1}^{N}\exp(\frac{p_{m}^{*}\cdot t_{j}}{\tau})})\\ &+\sum_{\{l|k_{l}=j\}}(1-\lambda_{l})\cdot\log(\frac{\exp(\frac{p_{l}^{*}\cdot t_{j}}{\tau})}{\sum_{m=1}^{N}\exp(\frac{p_{m}^{*}\cdot t_{j}}{\tau})})],\end{split}$ (13)
where $\tau$ represents the temperature hyperparameter, $\lambda$ denotes the voxel-mixing coefficient obtained from a Beta distribution where $\alpha=\beta=0.15$, and $N$ is the batch size. The CLIP embedding $t_{i}$ of a ground-truth image or text, denoted as $y_{i}$, is obtained as:
$t_{i}=\rm{CLIP}_{image/text}(y_{i}).$ (14)
The $p_{i}^{*}$ in Eq. (13) represents a fMRI embedding derived from either a low-level encoder $f_{L}$ or a high-level encoder $f_{H}$, denoted as $f_{L/H}$, using the initial fMRI $X_{i}$. This process is obtained with MixUp augmentation [71] as the subsequent equations:
$X_{\rm{mix}_{i,k_{i}}}=\lambda_{i}\cdot X_{i}+(1-\lambda_{i})\cdot X_{k_{i}},$ (15)
$p_{i}^{*}=f_{L/H}(X_{\rm{mix}_{i,k_{i}}}),$ (16)
After one third of the training stage, the BiMixCo loss for both $L_{image}$ and $L_{text}$ is replaced with soft contrastive loss to facilitate softmax probability distribution for improved knowledge distillation [73], expressed as:
$\mathcal{L}_{soft}=-\sum_{i=1}^{N}\sum_{j=1}^{N}[\frac{\exp(\frac{t_{i}\cdot t_{j}}{\tau})}{\sum_{m=1}^{N}\exp(\frac{t_{i}\cdot t_{m}}{\tau})}\cdot\log(\frac{\exp(\frac{p_{i}\cdot t_{j}}{\tau})}{\sum_{m=1}^{N}\exp(\frac{p_{i}\cdot t_{m}}{\tau})})].$ (17)
where the $p_{i}$ represents fMRI embeddings derived from either low-level or high-level encoder $f_{L/H}$ through the following equation:
$p_{i}=f_{L/H}(X_{i}).$ (18)
During inference, we exclusively use pre-trained weights for fMRI extraction $p_{i}=f_{L/H}(X_{i})$ and the subsequent 3D generation, relying entirely on fMRI data. This approach ensures that the model’s inference is based purely on its understanding of fMRI data, without requiring any information such as the viewed image, nor external texts.
4.6 3D generation evaluation
To analyze the generation quality, we introduce four fMRI-to-3D evaluation metrics: 3D Inception [36], 3D CLIP [37], 3D EfficientNet [38], and 3D Swav [39]. These metrics are designed to overcome the limitations of traditional pixel-by-pixel image reconstruction metrics like PixelCorrelation [76], which lacks robustness against rotations and translations in 3D space. For comprehensive evaluation, we construct a level-2 icosahedron with a 2.2 radius around the generated 3D scene and capture the 3D scene from all the icosahedron vertices to achieve multi-view coverage, similar to the approach in the text-to-3D benchmark [77]. The metrics 3D Inception and 3D CLIP, utilizing the final pooling layer of InceptionV3 [36] and CLIP model [37] respectively, assess the accuracy of two-way identification between original image embeddings and corresponding brain embeddings. Consequently, larger 3D Inception and 3D CLIP represent higher matching accuracy in features between ground truth images and generated 3D objects, indicating a well extracted features from fMRI. On the other hand, 3D EfficientNet and 3D Swav employ EfficientNet-B1 [38] and SwAV-ResNet50 [39] to calculate the average correlation distance between the extracted features. Lower correlation distance represents higher correlation between ground truth images and generated 3D objects, representing better reconstruction. In order to address view inconsistency in the final quality evaluation, we implement a regional convolution mechanism to smooth the scores across the local region, which treats the icosahedron as a graph composed of vertices and edges and performs mean pooling on this graph, employing a recursive formula:
$s_{i}^{(t+1)}=\frac{1}{\left|N(i)\right|+1}(s_{i}^{(t)}+\sum_{j\in N(i)}s_{j}^{(t)}).$ (19)
The highest score among all viewpoints is then selected as the ultimate quality score for the 3D generation.
4.7 Dataset
We conduct extensive experiments on the Natural Scenes Dataset (NSD) [29]. This dataset is comprised of high-resolution whole-brain fMRI measurements (1.8-mm isotropic resolution, 1.6-s sampling rate) from eight healthy adult subjects. These subjects were observed while viewing thousands of 2D colored natural images sampled from COCO dataset [74], spread across 30 to 40 scanning sessions. The two-stage training set includes 8,859 images and 24,980 fMRI trials, whereas the test set contains 982 images and 2,770 fMRI trials. To enhance data quality, we compute average fMRI responses for images that are presented more than once.
The fMRI signals are preprocessed and split into two specialized regions using predefined masks: the NSD General Region-of-Interest (ROI) mask and the Medial Temporal Lobe (MTL) mask, both at a resolution of 1.8 mm. The NSD General ROI mask encompasses regions such as V1, V2, V3, and V4, presenting a spectrum from the early visual cortex to higher visual areas, with an emphasis on low-level information. In contrast, the MTL mask encompasses brain regions including the hippocampus, amygdala, and parahippocampal regions, which are more attuned to high-level information.
4.8 Brain disorder diagnosis
Our model has the ability to translate complex fMRI signals into easily understandable multi-view object images. This empowers individuals, regardless of their professional background, to undertake disorder diagnosis with minimal training through a straightforward tutorial. This substantially lowers the barrier to entry for this task.
To simulate the disorders of varying brain regions, we adopt the general fMRI simulation techniques [75] that introduce Gaussian noises. We randomly select samples from the NSD test set for disorder simulation. For each sample, we apply independent Gaussian noises with a variance of 3.0 over five brain regions: V1, V2, V3, V4, and MTL, respectively.
In our diagnosis tutorial we show a participant several examples including the original image presented to a patient, and our model’s output under the healthy situation and regional disorder in V1, V2, V3, V4, and MTL, along with a textural description of each disorder’s characteristics. This tutorial can be easily conducted before the user study.
In our experiment, we invite 100 volunteer participants, each educated simply by the above tutorial. None of them possessed prior expertise in neuroscience or brain disorder diagnosis with fMRI. The diagnosis task is to detect which brain region has disorder by comparing the original image and the generated images by our model. For diagnosing a specific patient, we use a various number $\{1,5,10,20,50,100\}$ of images and fMRI pairs associated. We evaluate the diagnosis performance by the metrics of precision, recall, and F1 score.
5 Data Availability
Our research utilize the publicly available Natural Scenes Dataset (NSD) [29], which can be accessed freely for research and validation.
6 Code Availability
Our github repository is https://github.com/fudan-zvg/Brain3D. Supplementary materials with video visualizations of more 3D generations are available at https://brain-3d.github.io/.